OpenStudy (anonymous):

Which property of equality can be used to solve the equation n plus nine equals negative three? Addition Property of Equality Subtraction Property of Equality Multiplication Property of Equality Division Property of Equality

OpenStudy (anonymous):

I will give you a medal if you answer my question!

OpenStudy (anonymous):

k do you want help? @hotgurl72

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Sorry, I was locked out of my account, glitch.

OpenStudy (atlas):

n+9 = -3 Subtract 9 from both sides and they still remain equal n+9-9 = -3-9 n =-12

OpenStudy (atlas):

Which property do you think I have used

OpenStudy (anonymous):

Subtraction?

OpenStudy (atlas):

yes of course